OpenStudy (anonymous):

help please!! Find the zeroes of y = x(x − 4)(x + 2)

OpenStudy (anonymous):

Just set all of them to 0 x = 0 x-4= 0 x = 4 x+2 = 0 x = -2 x's are 0,4,-2

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

To find zero's of any x's you just have to set them all of them to 0 so you have 3 x's in that equation x (x-4) (x+2) so you set them all equal to 0 to find where the zeros exist thus you get x = 0 x-4=0 x+2=0 then you solve for x first one is already done x=0 second one x=4 third one x=-2 get it?

OpenStudy (anonymous):

yes thanks